Divide, and if possible, simplify. Assume all variables represent positie real numbers.
root:5(a^3b^9c^3)/sqrt:(a^-2bc)
I asked a tutor to assist and received a different answer than the possibilities that are on the quiz sheet!

Answer by RAY100(1637)

Please check to amke sure I am doing the right problem
.
(a^3b^9c^3)^(1/5) / (a^(-2) bc)^(1/2)
.
multiplying to eliminate powers,,,ie,,,(a^b)^c=a^(bc.
.
{a^(3/5) b^(9/5)c^(3/5) } / { a^(-1)b^(1/2)c^(1/2)}
.
now, term by term, subtract exponents,,ie,,a^b/a^c = a^(b-c)
.
a^{(3/5)-(-1)} b^{(9/5)/(1/2)} c^{(3/5)/(1/2)}
.
a^(8/5) * b^(13/10) *c^(1/10)
.
.
check: Let a=2,b=3,c=4,,both answer and original = 14.5,,,,ok